The frequency of a wave light is 1.0×10^6 sec^-.the wavelength for this wave is

Grade:12th pass

2 Answers

Arun
4 years ago
Wavelenght(lambda)=c/v=3×10^8/1×10^6=3×10^2m ans.
here c is the velocity of light and 
v is the frequency of given light
c is always constant for all electromagnetic radiations
Srija
4 years ago
  • V=c/lambda
  • Where c is nothing but velocity of light
  • v is frequency
  • Lambda is symbol of wave length
  • v= 3×108cm/1.0×106
  • v=3×10A